OpenStudy (anonymous):

please help!! 1 medal and fan If m

OpenStudy (anonymous):

42 + 42 = 84 m<C = m<A angles in a parallelogram = 360 2m<C = 360-84 = 276 m<C = 138

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you welcome. tnks for the medal!